# MobileVLM: A Fast, Strong and Open Vision Language Assistant for Mobile Devices

### v1

@@ 5.1 Ablation on Vision Backbones
| MobileLLaMA 1.4B | Alpaca | vicunav1 | 55.5 | 53.1 | 40.6 | 83.8 | 1168.0 | 47.7 |
| MobileLLaMA 1.4B | Vicuna | vicunav1 | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 |
@@ 5.4 Quantitative Analysis on SFT
Vicuna [132], fine-tuned on LLaMA, has been widely chosen in large multimodal models [75, 30, 133]. We further explore how SFT affects our language model’s performance in downstream tasks. Two common SFT paradigms Alpaca [111] and Vicuna [132] are compared in Table 10. We find that the scores of SQA, VQA, MME, and MMBench can all be significantly enhanced. It demonstrates that fine-tuning large language models in Vicuna dialogue mode [132] with the data from ShareGPT ultimately achieves the best performance. To better integrate SFT’s prompt format with the training of downstream tasks, we ablate the conversation mode on MobileVLM to find vicunav1 performs best.
@@ A.1 Dataset for MobileLLaMA
| Dataset | Proportion | |
| Arxiv | 2.5% | |
| Book | 4.5% | |
| C4 | 15.0% | |
| Common Crawl | 67.7% | |
| Github | 4.5% | |
| StackExchange | 2.0% | |
@@ Appendix B SFT prompt and response example
| Human: | who are you? |
| MobileLLaMA 1.4B Base: | I’m the AI. |
| MobileLLaMA 1.4B SFT: | I am a digital asset management and metadata expert. My expertise includes creating and managing digital content, metadata, and digital asset management solutions. I have a passion for helping organizations unlock the value of their digital assets and improve their workflows. I am also experienced in working with big data and analytics and can help organizations unlock insights from their data to drive better decision-making. I am always up for a challenge and love working with clients to help them achieve their goals. |
| Human: | what can you do? |
| MobileLLaMA 1.4B Base: | I’m a chatbot. |
| MobileLLaMA 1.4B SFT: | In addition to providing expert advice and guidance on a wide range of topics, I can also help you with tasks and projects that are more complex or require specialized knowledge, such as developing and implementing project plans or managing budget and resource allocation. Overall, my role is to provide value and support to your project or task and to help you achieve your goals in a timely and efficient manner. |
| System | A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. |
| USER | I want you to act as my travel agent for preparing an itinerary for travel to Munnar and Tekkady in Kerala. I have already booked flights from Hyderabad to Kochi for an onward journey on 25th Jan and a return journey on 28th Jan. We are a group of 4 men and planning to stay 2 days in Munnar and 1 day in Tekkady. I want you to help me with accommodation booking and activities. |

### v2

@@ 5.1 Ablation on Vision Backbones
| MobileLLaMA 1.4B | Alpaca | vicuna | 55.5 | 53.1 | 40.6 | 83.8 | 1168.0 | 47.7 |
| MobileLLaMA 1.4B | Vicuna | vicuna | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 |
@@ 5.4 Quantitative Analysis on SFT
Vicuna [132], fine-tuned on LLaMA, has been widely chosen in large multimodal models [75, 30, 133]. We further explore how SFT affects our language model’s performance in downstream tasks. Two common SFT paradigms Alpaca [111] and Vicuna [132] are compared in Table 10. We find that the scores of SQA, VQA, MME, and MMBench can all be significantly enhanced. It demonstrates that fine-tuning large language models in Vicuna dialogue mode [132] with the data from ShareGPT ultimately achieves the best performance. To better integrate SFT’s prompt format with the training of downstream tasks, we ablate the conversation mode on MobileVLM to find vicuna performs best.
@@ A.1 Dataset for MobileLLaMA
| Dataset | Proportion |
| Arxiv | 2.5% |
| Book | 4.5% |
| C4 | 15.0% |
| Common Crawl | 67.7% |
| Github | 4.5% |
| StackExchange | 2.0% |
@@ Appendix B SFT prompt and response example
Human: |
who are you? |
MobileLLaMA 1.4B Base: |
I’m the AI. |
MobileLLaMA 1.4B SFT: |
I am a digital asset management and metadata expert. My expertise includes creating and managing digital content, metadata, and digital asset management solutions. I have a passion for helping organizations unlock the value of their digital assets and improve their workflows. I am also experienced in working with big data and analytics and can help organizations unlock insights from their data to drive better decision-making. I am always up for a challenge and love working with clients to help them achieve their goals. |
Human: |
what can you do? |
MobileLLaMA 1.4B Base: |
I’m a chatbot. |
MobileLLaMA 1.4B SFT: |
In addition to providing expert advice and guidance on a wide range of topics, I can also help you with tasks and projects that are more complex or require specialized knowledge, such as developing and implementing project plans or managing budget and resource allocation. Overall, my role is to provide value and support to your project or task and to help you achieve your goals in a timely and efficient manner. |
System |
A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. |
USER |
I want you to act as my travel agent for preparing an itinerary for travel to Munnar and Tekkady in Kerala. I have already booked flights from Hyderabad to Kochi for an onward journey on 25th Jan and a return journey on 28th Jan. We are a group of 4 men and planning to stay 2 days in Munnar and 1 day in Tekkady. I want you to help me with accommodation booking and activities. |
ASSISTANT |